# Continual-MEGA: A Large-scale Benchmark for Generalizable Continual Anomaly Detection

### v2

@@ 3 Proposed Baseline Method
(a) Training scenario of $Base$ and Task $N$.
(b) Inference process. $B$ and $N$ denote the adapters corresponding to the base classes and the task-specific classes $\{1,2,3,...,N\}$, respectively. To inference, we use $A^{\text{avg}}$, which is the average of the adapter weights trained on the base classes and the $N$ task adapters.
Figure 4: Overview architecture.
3.1 Mixture-of-Expert of Adapters
To solve the Continual-MEGA benchmark, we propose a new baseline AD method, using CLIP radford2021learning text and visual encoders. The diagrams in Figures 4(a) and 4(b) represent the overall framework of the method. We use a set of four-adapter $A=\{A_{1},\dots A_{4}\}$ for each block of layers $S_{1}$ to $S_{4}$ of the CLIP visual encoder. For each categories including the set of (Base) classes $C_{b}$ and the set of $n$’th task $C_{n}$ classes, where the entire task number is $n=1,\dots N$, we separately train the adapter set denoted as $A_{n}=\{A_{1,B},\dots,A_{4,B}\}$. The features $F_{1}$ to $F_{4}$ from the adapters are added to form features to represent the anomaly map, and we apply the proposed loss function for training all the adapters. Specifically, we use the Text encoder converted features for preset normal and abnormal text prompts to supplement AD training. The text features $F_{\text{text}}\in\mathbb{R}^{2\times d}$ for normal and anomaly are obtained from a text encoder through the text prompt $p$. The detailed settings for the text prompt are available in the appendix A.3.
In the inference stage, we accumulate all the pre-trained adapters $A_{n}=\{A_{1,B},\dots,A_{4,B}\}$ and $A_{b}$ for each task and baseline class set, by the average adapters $A^{\text{avg}}=\{A^{\text{avg}}_{1},...,A^{\text{avg}}_{4}\}$, as follows:
$A^{\text{avg}}_{l}=\frac{1}{N+1}(\sum^{N}_{n}{A_{l,n}}+A_{l,b}).\end{aligned}$ (1)
where the number $l=\{1,\dots 4\}$ denotes the ordering of each of four blocks. In implementation, each adaptation layer $A_{l}(\cdot)$ consists of two linear layers as:
$A_{l}(F_{l})=W_{l,2}(W_{l,1}F_{l}^{T}),$ (2)
where $F_{l}\in\mathbb{R}^{G\times d}$ represents the visual features extracted from the visual encoder stage $S_{l}$, whereas $G$ and $d$ present the number of grids and feature dimension of each grid, respectively. $W_{l,1}$ and $W_{l,2}$ denote the learnable parameters of linear transformations. To preserve the original knowledge from the pre-trained CLIP model, the input $\tilde{F}_{l}$ to $S_{l+1}$ is designed by a residual connection (gao2024clip, ) as
$\tilde{F}_{l}=\alpha F_{l}+(1-\alpha)A_{l}(F_{l}),$ (3)
where $\alpha$ represents a residual ratio between the original features and the adapted features. $\alpha$ is set to 0.9 in our work. We use the CLIP with ViT-L/14 dosovitskiy2020image architecture, which consists of 24 sublayers divided into four layers, where each layer contains six sublayers. The size of input images was set to 336. The adaptation layers for anomaly feature generation were applied to layers 1, 2, 3, and 4.
3.2 Synthetic Feature Generation
Specifically, we apply random noise to enable the adaptation layers $A_{l}$ to learn a diverse range of anomalies. In training, we use task-wise zdators $A_{n}$ and in the inference phase, we use the accumulated adators $A_{avg}$. The synthetic anomaly features ($F_{l}^{1}$) are generated by
$F_{l}^{1}=A_{l}(F_{l}+\gamma),$ (4)
where $\gamma\in\mathbb{R}^{G\times d}$ is a random noise. The adapted normal features ($F_{l}^{0}$) are generated via the adaptation layers as
$F_{l}^{0}=A_{l}(F_{l}).$ (5)
The adapted normal features $F_{l}^{0}$ and synthetic anomaly features $F_{l}^{1}$ are booth used to generate anomaly score maps through cosine similarity operations along with the text features.
3.3 Model Training
The proposed loss function $L(\cdot)$ are implemented to detect pixel-wise anomalies. For each layer $l$, we define pixel-wise losses $L_{\text{no}}$, $L_{\text{an}}$, $L_{\text{syn}}$, each representing the losses for normal, anomalies, and synthetic features, as:
$\mathcal{L}_{\text{no}}=L_{\text{ce}}+L_{d}+L_{f},$ (6)
$\mathcal{L}_{\text{an}}=\mathcal{L}_{d}+L_{f},$ (7)
$\mathcal{L}_{\text{syn}}=L_{\text{ce}}.$ (8)
The losses termed as $L_{\text{ce}}$, $L_{f}$ and $L_{d}$ are the pixel-level cross-entropy, focal (ross2017focal, ) and dice  (milletari2016v, ) loss. All three losses get features $F_{l}$ as input, and get feature mask $M_{l}$ as label. For the mask $M_{l}$, the real-anomaly region and synthetic feature region are masked with ones, and the real-normal region is masked as 0. Finally, the total loss for the layer $l$ is computed as the summation of the previously defined losses in equation (6), (7) and (8), as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{no}}+\mathcal{L}_{\text{an}}+\mathcal{L}_{\text{syn}}.$ (9)
For training, we accumulate all the layerwise loss functions to update adapters.
4 Related Works
Anomaly detection is a specialized task focused on detecting and rejecting unknown samples amodei2016concrete ; hendrycks2021unsolved , often framed as an out-of-distribution (OOD) problem, particularly targeting industrial anomalous data chalapathy2019deep ; pang2021deep ; ruff2021unifying . Specifically, anomaly detection involves both identifying image-level anomalies and localizing defective regions. However, object categories exhibit diverse characteristics, and the detection challenges vary significantly across these categories. Due to the nature, early deep anomaly detection models suppose per-class, or unsupervised anomaly detection bergmann2022beyond ; cohen2020sub ; defard2021padim ; li2021cutpaste ; ristea2022self ; roth2022towards ; zavrtanik2021draem ; zou2022spot .
Toward Unified Anomaly Detection.
Recent advances in large-scale backbone models, such as CLIP radford2021learning , offer a promising solution to the challenge of unified anomaly detection across categories jeong2023winclip ; he2024learning ; yao2024hierarchical ; strater2024generalad . Following the initial approach jeong2023winclip utilizing CLIP, current trends are focused on developing unified anomaly detection models with zero- and few-shot adaptation capabilities across categories.
Zero- and Few-shot Adaptation.
Anomaly detection with zero- and few-shot adaptation jeong2023winclip ; li2024promptad ; chen2023zero ; tamura2023random ; gu2024filo ; gui2024few ; qu2024vcp ; zhou2023anomalyclip ; deng2023anovl ; chen2024clip across various categories reflects real-world scenarios where acquiring a sufficient number of samples for newly incoming categories is often not feasible, and obtaining anomaly samples is even more challenging. In the few-shot adaptation scenario, typically one to five normal or anomaly samples are used to fine-tune the original adaptation models. The zero-shot adaptation scenario assumes no adaptation across categories. Various methods have been proposed to address these challenges, including text prompt utilization jeong2023winclip ; li2024promptad ; zhou2023anomalyclip ; deng2023anovl ; chen2024clip ; tamura2023random ; gu2024filo , visual context prompting qu2024vcp ; deng2023anovl , and anomaly dataset synthesis chen2023zero ; chen2024unified . Notably, most of these approaches leverage text prompt information, with strategies ranging from manually designed templates jeong2023winclip ; deng2023anovl , normal sample-only li2024promptad , learned zhou2023anomalyclip ; deng2023anovl ; gu2024filo , to augmented prompting tamura2023random .
Continual Adaptation.
Another important aspect in recent anomaly detection research trend is continual adaptation li2022towards ; pang2025context ; liu2024unsupervised ; tang2024incremental ; jin2024oner ; meng2024moead ; mcintosh2024unsupervised , which mirrors the scenario where object categories arrive incrementally. In this context, we aim to mitigate catastrophic forgetting while ensuring that adaptation to previous categories improves the model’s performance for future category adaptations. Building on initial efforts li2022towards , several approaches have been proposed, including context-aware feature adaptation pang2025context , learned text prompts liu2024unsupervised , integration of a unified reconstruction-based detection framework tang2024incremental , online replay memory design jin2024oner , parameter-efficient tuning meng2024moead , and unsupervised tuning mcintosh2024unsupervised ; tang2024incremental . The continual evaluation scenario is built on widely-used public datasets such as MVTec-AD bergmann2019mvtec or VisA zou2022spot , but the quantity and diversity of the dataset are limited compared to the continual adaptation scenario bang2021rainbow using ImageNet deng2009imagenet .
Building upon existing studies, the core novelty of our work lies in introducing a new research agenda for continual anomaly detection. First, we propose Continual-MEGA, a large-scale evaluation benchmark designed to measure overall AD performance in more challenging and scalable scenarios. Second, we present a novel AD method that integrates efficient CLIP-based adaptation with anomaly feature synthesis and optimized prompt tuning. Unlike prior approaches such as chen2024unified ; chen2023zero ; meng2024moead , our method demonstrates robust performance across the newly proposed benchmark, highlighting its effectiveness under more demanding evaluation settings.
5 Experiments
Overview.
Tables 1 and 2 present the quantitative results under the proposed evaluation scenarios, comparing various recently proposed anomaly detection methods. In our experiments, we categorize the methods into three groups: (1) approaches that adapt using only normal samples: SimpleNet liu2023simplenet , GeneralAD strater2024generalad , HGAD yao2024hierarchical , and ResAD yao2024resad , (2) vision-language model (VLM)-based methods: MVFA huang2024adapting , VCP-CLIP qu2024vcp , and MediCLIP zhang2024mediclip , and (3) methods specifically designed for continual learning settings: UCAD liu2024unsupervised , and IUF tang2024incremental . Overall, the results indicate a substantial drop in performance across all methods—particularly for pixel-wise anomaly detection—when evaluated under the proposed continual learning settings. This contrasts sharply with the higher performance typically observed in standard benchmarks such as MVTec-AD and VisA, highlighting the increased difficulty and practical relevance of our evaluation protocol.
Evaluation of Continual Learning Capability.
Scenario 1 represents a typical continual learning setup but significantly scales up both the number of classes and data volume compared to prior works on continual adaptation liu2024unsupervised ; tang2024incremental . Notably, vision-language model (VLM)-based methods such as MVFA zhang2024mediclip and MediCLIP zhang2024mediclip achieve the highest performance among all baselines, apart from the proposed method. Specifically, MVFA and our proposed method achieved comparable performance to each other. The overall results strongly suggest that the existing methods, including continual learning approaches for anomaly detection, struggle to handle large-scale continual evaluation settings.
Table 1: Experimental results on Scenario 1. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. While all methods were trained with the same number of epochs for fair comparison, the $\text{IUF}^{\ast}$ method requires substantially longer training due to its methodology. Therefore, we trained the $Base$ classes for 500 epochs and the $New$ classes for 100 epochs in the $\text{IUF}^{\ast}$ setting. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
@@ 5 Experiments
This observation reveals two key insights: (1) several prior methods appear to be tightly fitted to existing benchmarks such as MVTec-AD and VisA, which limits their generalizability to more diverse or challenging settings, and (2) methods with stronger initial (pretrained) performance tend to retain higher accuracy throughout continual adaptation. Regarding the first insight, methods such as MVFA, which demonstrate competitive performance in Scenario 1, exhibit significantly degraded results, particularly in pixel-level AP, when MVTec-AD and VisA datasets are excluded from the Base classes and New classes, as shown in Table 2. In contrast, our proposed method consistently achieves robust performance across all evaluation scenarios. In our proposed setup, detecting the baseline categories is substantially more challenging than in conventional benchmarks, as shown in Table 3. Under this setting, VLM-based methods demonstrate significantly stronger performance compared to methods explicitly designed for continual learning. This discrepancy can be attributed to the limited detection capability of existing continual anomaly detection methods, even at their initial stage, a factor closely tied to the second insight discussed earlier. Consequently, these methods face greater difficulty in adapting to new incoming categories. Although the forgetting measure (FM) of continual learning-based methods appears lower than that of VLM-based methods, this is likely due to their poor initial detection performance rather than effective forgetting mitigation.
@@
Table 2: Experimental results on Scenario 2. To evaluate the zero-shot generalization performance of the methods, we excluded the MVTec-AD and VisA classes from training and used them only for evaluation. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. Zero-shot performance on MVTec-AD and VisA is presented in Figure 5. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
  Type Method 58-5 (12 tasks) 58-10 (6 tasks) 58-30 (2 tasks)
Only-normal SimpleNet 56.1/4.2/30.2 8.2/1.4/4.8 56.5/3.8/30.2 7.1/2.4/4.8 57.3/3.8/30.6 4.9/1.0/3.0
GeneralAD 49.0/0.8/24.9 6.3/1.7/4.0 51.3/0.9/26.1 3.1/1.2/2.2 47.7/2.1/24.9 5.7/0.0/2.9
HGAD 51.1/4.3/27.7 1.8/0.3/1.1 51.8/4.5/28.2 1.4/0.2/0.8 51.8/4.3/28.1 2.4/0.2/1.3
ResAD 48.8/0.6/24.7 10.7/1.0/5.8 42.7/1.7/22.2 3.0/0.9/1.9 55.6/12.8/34.2 12.0/4.7/8.3
VLM-based MVFA 63.2/4.7/34.0 5.8/5.4/5.6 64.0/4.1/34.1 5.8/2.9/4.4 65.3/5.0/35.2 1.9/2.0/2.0
VCP-CLIP 55.6/18.7/37.1 3.8/6.8/5.3 53.2/19.8/36.5 0.3/3.0/1.7 64.3/22.3/48.3 2.8/3.7/3.3
MediCLIP 79.6/7.3/43.5 3.8/5.6/4.7 76.0/6.0/41.0 4.9/3.7/4.3 77.1/5.9/41.5 2.1/7.0/4.6
Continual UCAD 66.0/7.4/36.7 0.4/0.02/0.2 63.5/6.0/34.8 0.7/0.03/0.4 58.0/3.1/30.6 0.0/0.0/0.0
IUF 57.6/4.2/30.9 1.7/0.5/1.1 58.0/4.3/31.2 0.3/0.2/0.3 58.0/4.3/31.2 -0.7/-0.1/-0.4
$\text{IUF}^{\ast}$ 60.2/6.3/33.3 0.8/0.3/0.6 60.7/6.4/33.6 0.2/0.1/0.2 61.7/7.0/34.4 0.2/0.2/0.2
Ours 71.7/20.7/46.2 2.3/4.1/3.2 72.4/22.2/47.3 2.5/3.8/3.2 76.8/27.5/52.2 1.0/2.6/1.8
Evaluation of Generalizability after Continual Adaptation.
Another notable contribution of our proposed evaluation protocol is the incorporation of zero-shot generalization evaluation following continual adaptation, defined as Scenario 2 in Table 2. This setting reflects practical situations where models must retain or even improve zero-shot capability during continual learning. However, this vital aspect has been largely overlooked in previous anomaly detection benchmarks, which do not adequately capture the demands of real-world deployment. As shown in Table 2 and the accompanying visualizations in Figure 5, only a few methods, most notably MediCLIP, demonstrate decent generalizability after adaptation. We conjecture that meaningful analysis of zero-shot trends is only possible for methods that already exhibit sufficient generalization capability prior to adaptation. Specifically, by comparing results with Scenario 3 in Figure 5, where the proposed ContinualAD dataset is excluded from the Base classes and New classes, we observe that only methods with robust performance in continual and zero-shot generalization settings exhibit performance gains when additional categories are introduced. The full comparison results for the continual learning setting in Scenario 3 are provided in the appendix B.
Summarization.
The overall quantitative results highlight three key takeaways. First, under our expanded evaluation setup, most existing methods, including those explicitly designed for continual learning, exhibit insufficient detection performance, limiting their utility for meaningful analysis. Specifically, the pixel-level AP reported under our proposed evaluation setting shows a substantial performance gap compared to existing benchmarks, with significantly lower scores across most methods. Second, a small subset of methods, primarily VLM-based approaches, demonstrate relatively strong performance and benefit from decent generalization after continual adaptation. From this perspective, Scenario 3 further supports the value of our proposed ContinualAD dataset, showing that it provides meaningful supervision signals that improve detection capability for future incoming categories. Most importantly, the proposed baseline method consistently outperforms existing approaches across all scenarios, establishing a strong benchmark and emphasizing the significance of both the proposed method and the newly introduced continual evaluation protocol.
(a) Scenario 2.
(b) Scenario 3.
Figure 5: Zero-shot performance comparison on MVTec-AD and VisA under Scenario 2 and Scenario 3. The upper two plots present results from Scenario 2, while the remaining plots correspond to Scenario 3. Each point denotes the performance of an AD method, with image-level AUROC on the x-axis and pixel-level AP on the y-axis. Best viewed in wide.
Table 3: Experimental results of base classes across scenarios. We note that the base classes used in Scenario 2 and Scenario 3 differ, as ContinualAD is included among the base classes in Scenario 2 but not in Scenario 3. The best-performing results are highlighted in bold.
Image Pixel Image Pixel Image Pixel
Only-normal SimpleNet 58.8 6.3 61.3 4.5 57.5 4.5
GeneralAD 51.5 2.6 52.6 1.8 54.4 2.7
HGAD 59.5 5.0 56.1 3.2 55.5 2.7
ResAD 73.3 15.5 69.1 7.8 70.7 15.3
VLM-based MVFA 81.7 32.6 65.8 10.4 70.7 21.2
VCP-CLIP 73.8 25.4 61.0 23.1 61.9 22.5
MediCLIP 73.9 4.5 78.1 8.5 75.3 5.9
Continual UCAD 55.8 1.6 58.1 4.7 56.0 3.6
IUF 60.5 7.4 57.4 4.4 58.5 4.2
$\text{IUF}^{\ast}$ 68.3 13.5 65.8 9.5 63.6 9.5
Ours 83.1 39.0 82.0 35.7 77.8 36.5

### v3

@@ 4 Proposed Baseline Method
To address the Continual-MEGA benchmark, we propose a simple baseline AD method, Anomaly Detection across Continual Tasks (ADCT), which leverages the text and visual encoders of CLIP (Radford et al., 2021), drawing inspiration from AnomalyCLIP (Deng and Li, 2022). ADCT is designed to fully leverage CLIP information with minimal modification. ADCT employs a lightweight MLP adapter for feature adaptation in each CLIP block, in conjunction with a feature synthesis module, design choices that may contribute to effective CZSL performance while preserving pre-trained CLIP knowledge. Figure 4 illustrates the incremental training process on the $Base$ classes and successive Task $N$ classes, while Figure B in the Appendix depicts the inference pathway formed by a mixture-of-experts adapters.
@@
Table 1: Experimental results on Scenario 1. $\cdot/\cdot/\cdot$ denotes the Image-AUROC, Pixel-AP and their average, respectively. While all methods were trained with the same number of epochs for fair comparison, the $\text{IUF}^{\ast}$ method requires substantially longer training due to its methodology. Thus, we trained the $Base$ classes for 500 epochs and the $New$ classes for 100 epochs in the $\text{IUF}^{\ast}$ setting. The notation “$X$-$Y$ ($Z$ tasks)” in the first row denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Z$ continual learning phases (tasks), each comprising $Y$ new classes.
4.1 Mixture-of-Experts of Adapters
To design the adapter, we employ a set of four adapters $A=\{A_{1},\dots A_{4}\}$ for each block of layers of the CLIP visual encoder. For each category including the set of (Base) classes $C_{b}$ and the set of $n$’th task $C_{n}$ classes, where the total number of tasks is $n=1,\dots N$, we separately train the adapter set denoted as $A_{n}=\{A_{1,n},\dots,A_{4,n}\}$.
In the inference stage, as illustrated in Appendix Figure B, we accumulate all the pre-trained adapters $A_{n}$ and $A_{b}$ for each task and Base classes set, by the average adapters $A^{\text{avg}}=\{A^{\text{avg}}_{1},...,A^{\text{avg}}_{4}\}$, as follows:
$\begin{aligned} A^{\text{avg}}_{l}=\frac{1}{N+1}(\sum^{N}_{n}{A_{l,n}}+A_{l,b}).\end{aligned}$ (1)
where the number $l=\{1,\dots 4\}$ denotes the index of each of the four blocks. In implementation, each adaptation layer $A_{l}(\cdot)$ consists of two linear layers as:
$A_{l}(F_{l})=W_{l,2}(W_{l,1}F_{l}^{T}),$ (2)
where $F_{l}$ represents the visual features extracted from the $l$-th visual encoder stage of CLIP. We use CLIP with ViT-L/14 (Dosovitskiy et al., 2020) architecture, which consists of 24 sublayers divided into four layers, where each layer contains six sublayers. The size of input images was set to 336. The adaptation layers for anomaly feature generation were applied to layers 1, 2, 3, and 4.
4.2 Synthetic Feature Generation
Specifically, we apply random noise to enable the adaptation layers $A_{l}$ to learn a diverse range of anomalies. In training, we use task-wise adapters $A_{n}$ and in the inference phase, we use the accumulated adapter $A^{\text{avg}}$. The synthetic anomaly features ($F_{l}^{1}$) are generated by
$F_{l}^{1}=A_{l}(F_{l}+\gamma),$ (3)
where $\gamma\in\mathbb{R}^{G\times d}$ is random noise. The adapted normal features ($F_{l}^{0}$) are generated via the adaptation layers as
$F_{l}^{0}=A_{l}(F_{l}).$ (4)
The adapted normal features $F_{l}^{0}$ and synthetic anomaly features $F_{l}^{1}$ are both used to generate anomaly score maps by calculating cosine similarity along with the text features.
5 Experiments
Table 2: Experimental results on Scenario 3. To verify the effectiveness of the ContinualAD dataset, it is excluded from the training process. $\cdot/\cdot/\cdot$ denotes the Image-AUROC, Pixel-AP, and their average, respectively. The notation “$X$-$Y$ ($Z$ tasks)” denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Z$ continual learning phases (tasks), each comprising $Y$ new classes.
Overview.
Tables 1 and 2 present the quantitative results under the proposed evaluation scenarios, comparing various recently proposed anomaly detection methods. In our experiments, we categorize the methods into three groups: (1) approaches that adapt using only normal samples: SimpleNet (Liu et al., 2023), GeneralAD (Sträter et al., 2024), HGAD (Yao et al., 2024a), and ResAD (Yao et al., 2024b); (2) vision-language model (VLM)-based methods: MVFA (Huang et al., 2024), VCP-CLIP (Qu et al., 2024), and MediCLIP (Zhang et al., 2024a); and (3) methods specifically designed for continual learning settings: UCAD (Liu et al., 2024) and IUF (Tang et al., 2024). Overall, the results indicate a substantial drop in performance across all methods, particularly for pixel-wise anomaly detection, when evaluated under the proposed continual learning settings. This contrasts sharply with the higher performance typically observed in standard benchmarks such as MVTec-AD and VisA, highlighting both the increased difficulty and practical relevance of our evaluation.
Table 3: Experimental results on Scenario 3. To verify the effectiveness of the ContinualAD dataset, it is excluded from the training process. $\cdot/\cdot/\cdot$ denotes the Image-AUROC, Pixel-AP, and their average, respectively. The notation “$X$-$Y$ ($Z$ tasks)” denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Z$ continual learning phases (tasks), each comprising $Y$ new classes.
Figure 5: Learning curves on Scenario 2 (10 classes per task). We plot the average of image-level AUROC and pixel-level AP over all tasks seen so far as the model is incrementally trained from Task 1 to Task 6. This plot highlights that under matched compute, ADCT (Ours) exhibits superior stability, preserving performance across tasks substantially better than MVFA, which corroborates our quantitative FM results in Table 2.
       Components           58-30 (2 tasks)           Zero-shot
Adapters Synthetic Mixture Image Pixel MVTec-AD VisA
ACC FM ACC FM Image Pixel Image Pixel
56.0 1.0 75.2 2.3 61.8 1.0
$77.7\pm 0.6$ $0.6\pm 0.4$ $22.3\pm 3.1$ $0.3\pm 0.3$ $86.9\pm 1.0$ $26.3\pm 4.0$ $82.8\pm 0.1$ $13.9\pm 3.8$
$77.2\pm 0.5$ $4.1\pm 0.2$ $30.1\pm 1.1$ $6.7\pm 1.7$ $82.5\pm 1.2$ $35.7\pm 0.2$ $77.7\pm 3.3$ $19.7\pm 0.7$
$76.3\pm 0.4$ $1.4\pm 0.5$ $26.8\pm 0.8$ $2.8\pm 0.2$ $81.2\pm 0.8$ $32.1\pm 0.8$ $78.8\pm 0.2$ $18.8\pm 0.3$
Table 4: Ablation study on Scenario 2. All results except the first row are averaged over three random seeds and reported as mean $\pm$ standard deviation. The first row without any components corresponds to vanilla pretrained CLIP without any continual learning; thus FM is not defined (shown as “–”), and the reported scores are single values rather than mean $\pm$ standard deviation.
Evaluation of Continual Learning Capability.
Scenario 1 represents a typical continual learning setup but significantly scales up both the number of classes and the volume of data compared to prior works on continual adaptation (Liu et al., 2024; Tang et al., 2024).
Notably, vision-language model (VLM)-based methods such as MVFA (Huang et al., 2024) and MediCLIP (Zhang et al., 2024a) achieve the highest performance among all baselines, apart from the proposed method. Specifically, MVFA and our proposed method show comparable performance. The overall results strongly suggest that existing methods, including continual learning approaches for anomaly detection, struggle to handle large-scale continual evaluation settings. This observation reveals two key insights: (1) several prior methods appear to be tightly fitted to existing benchmarks such as MVTec-AD and VisA, which limits their generalizability to more diverse or challenging settings; and (2) methods with stronger initial (pretrained) performance tend to retain higher accuracy throughout continual adaptation.
Regarding the first insight, methods such as MVFA, which demonstrate competitive performance in Scenario 1, exhibit significantly degraded results, particularly in pixel-level AP, when MVTec-AD and VisA datasets are excluded from the Base and New classes, as shown in Table 2. In contrast, our proposed method consistently achieves robust performance across all evaluation scenarios. Regarding the second insight, VLM-based methods demonstrate significantly stronger performance compared to methods explicitly designed for continual learning. This discrepancy can be attributed to the limited detection capability of existing continual anomaly detection methods, even at their initial stage. Consequently, these methods face greater difficulty in adapting to new incoming categories. Although the forgetting measure (FM) of continual learning-based methods appears lower than that of VLM-based methods, this is likely due to their poor initial detection performance rather than effective forgetting mitigation.
Evaluation of Generalizability and Stability.
Another notable aspect of the evaluation is the CZSL scenario, which evaluates zero-shot generalization following continual adaptation. In Scenario 2, our proposed baseline model demonstrates improved generalization, benefiting from both a stronger set of Base classes and the continual adaptation of additional classes. Scenario 2 includes our proposed ContinualAD dataset, under which most methods exhibit improved zero-shot performance after continual learning. In contrast, when ContinualAD is excluded in Scenario 3, most existing methods suffer a degradation in zero-shot generalization after continual learning. Among them, our proposed method shows the smallest performance drop, indicating stronger robustness to continual adaptation compared to other approaches. To further investigate this observation, we refer to the quantitative results from Scenarios 2 and 3, presented in Table 2 and Table 3. These tables provide detailed evaluation metrics corresponding to continual adaptation and zero-shot evaluation results. To facilitate comparison, we visualize the results for representative methods of the tables in Figure C of the Appendix.
To further analyze how performance evolves as new tasks arrive, Figure 5 presents the learning curves in Scenario 2 with 10 classes per task and six tasks in total. For each increment (Task 1 $\rightarrow$ Task 6), we plot the average of image-level AUROC and pixel-level AP over all tasks observed so far, under a matched-compute setting (identical epochs, input resolution, and continual stream). Across all increments, MVFA consistently underperforms ADCT and exhibits a sharper degradation as new tasks are introduced, indicating more pronounced forgetting of earlier tasks, whereas ADCT maintains a higher and flatter curve, demonstrating superior stability. These observations are consistent with our FM results and underscore that explicitly controlling forgetting is crucial for stable continual anomaly detection.
Notably, we also observe that methods such as MediCLIP (Zhang et al., 2024a) and our proposed ADCT, both leveraging CLIP with lightweight adapters and image or feature synthesis, consistently achieve competitive performance in the CZSL scenario. Based on these results, we can conjecture that our ContinualAD dataset provides valuable information for detecting anomalies in both unseen and continually introduced categories under more challenging scenarios than prior setups. Our baseline demonstrates robust and consistent performance, suggesting a potentially valuable insight: lightweight CLIP adaptation combined with feature synthesis may contribute to improved continual and CZSL learning.
Ablation Study.
Table 4 presents an ablation on Scenario 2 (30 classes per task), isolating three components: the adapter-based continual learner (Adapters), synthetic anomaly generation (Synthetic), and the mixture-of-expert adapters inference scheme (Mixture). The first row corresponds to vanilla CLIP without any of these components, so FM in the continual setting is undefined and omitted. All configurations with at least one component outperform vanilla CLIP in both the continual setting and the zero-shot evaluations. Comparing rows with and without Synthetic, we observe that omitting synthetic anomaly generation (Adapters + Mixture) leads to noticeably lower pixel-level scores in both the continual setting and the zero-shot evaluation, highlighting that synthetic anomalies are crucial for strong pixel-level performance. Since image-level metrics can be overly optimistic when anomaly localization is poor, we view pixel-level scores as a more reliable indicator of anomaly detection quality. Using Adapters with Synthetic but without Mixture (Adapters + Synthetic) gives high overall performance but substantially larger FM, indicating strong forgetting of previous tasks. In contrast, the full configuration (Adapters + Synthetic + Mixture) maintains strong zero-shot performance while significantly reducing FM, which is desirable for a continual-learning baseline that aims to minimize forgetting without sacrificing zero-shot accuracy.